A square sheat of steal has a side of 15cm at 00 C. Determine its area at 400C. (Linear expansivity of steel = 1.1x10-5 0C-1
1
Expert's answer
2020-06-03T12:14:00-0400

Length of the square plate is at "T_i=0^{\circ}C" is "L_0=15cm = 0.15m" ,thus "A_0=L_0^2=0.0225m^2" ,"T_f=400^{\circ}C"

"\alpha_L=1.1\times 10^{-5}K^{-1}" .

Let,"A" the required area and "\Delta T=T_f-T_i=400K" .


We know that ,

"A=A_0(1+2\alpha_L \Delta T)"

Thus on plugin the values in above equation we get,

"A=0.0225(1+2\times 1.1\times 10^{-5} \times400)\\
\implies A=0.022698m^2"